# Inflated hot Jupiters: inferring average atmospheric velocity via Ohmic models coupled with internal dynamo evolution

### v1

@@ 4.2 Atmospherically induced current
Our aim is to estimate the spherically averaged radial profile of the current density, $J(r,t):=<|\@vec{J}|>_{\theta,\phi}(r,t)$ (where ${J}=(\@vec{\nabla}\times\@vec{B})/\mu_{0}$, and $\mu_{0}$ is the magnetic permeability), in order to evaluate the radial profile of the local Joule dissipation rate, $Q_{j}(r,t)=J(r,t)^{2}/\sigma(r,t)$, in the 1D code. As in the first works on Ohmic dissipation in HJs (Batygin & Stevenson, 2010; Batygin et al., 2011), we can quantify the electrical currents by imposing a stationary solution to the induction equation, given by the balance between the advection (itself given in terms of the thermal wind ${v}$), and the resistive term (neglecting the Hall and ambipolar non-ideal terms, see Perna et al. 2010a; Koskinen et al. 2014; Soriano-Guerrero et al. 2025):
$\frac{d\@vec{B}_{\rm ind}}{dt}=\@vec{\nabla}\times\left(\@vec{v}\times\@vec{B}-\frac{\@vec{J}}{\sigma}\right)=0\leavevmode\nobreak\ ,$ (7)
where the magnetic field ${B}$ includes the dynamo-generated background field, ${B}_{\rm dyn}$, and the atmospherically induced field itself, ${B}_{\rm ind}$. Following e.g. Batygin & Stevenson (2010), at equilibrium, the solution of the induced current density can be then defined by considering the electrical potential $\Phi$:
@@
and imposing the continuity equation ${J}=0$, so that
@@
We can then expand $\Phi$ in spherical harmonics, $\Phi=\Sigma_{lm}\Phi_{lm}(r)Y_{l}^{m}(\theta,\phi)$, and solve for the radial dependence of each multipole $\Phi_{lm}(r)$, for a given prescription of ${v}$, ${B}$ and $\sigma$ (which in turn depends on $T$ and density, i.e., on $p(T)$). Batygin & Stevenson (2010), Batygin et al. (2011) and related works solve such elliptic equation, considering first the uppermost regions where ${v}\neq 0$, and assuming axial symmetry (i.e., $\Phi_{lm}=0$ for any $m\neq 0$) and simplified geometries for ${v}(r,\theta)$, ${B}(r,\theta)$, for a given profile $\sigma(r)$. The simplest configuration is a purely zonal wind $v_{\phi}\sim\sin\theta$ ($\theta\in[0,\pi]$ being the co-latitude) and a dipolar, aligned field, which means that the only possible induced current multipole is the quadrupolar one, $\Phi_{20}\neq 0$.
Here we propose new families of semi-analytical solutions for the deep layers where ${v}\sim 0$, treating the complex atmospheric layers in an effective, parametrized way, and allowing for any generic profile $\sigma(r)$.
Figure 2: Electrical conductivity $\sigma$ as a function of pressure for different planetary models. We show ages of $t\simeq 0.4$ (dots), 1.1 (dashes), and 5 Gyr (solid lines), for a planet with $M=1\leavevmode\nobreak\ M_{j}$, showing two cases $T_{\rm eq}=1500,2000$ K, without (black, grey) and with Ohmic heating, parameter $v_{\rm avg}=3000$ and 356 m/s, respectively (blue, red). We also show (brown and gold lines), for comparison, the simplified $K^{+}$-only thermal ionization formula, which neglects pressure ionization, for the cases without Ohmic heating.
Figure 3: Profiles as a function of pressure of (from top to bottom): induced currents $J$ from eq. (15), Ohmic specific and volumetric heat rate, for the same cases and ages as Fig. 2. The thick lines in the bottom panel identify the convective regions at each age, of each model.
@@ 4.2.1 Treatment of the induction in the atmospheric region
There are several factors that introduce uncertainties and, likely, planet-to-planet variations in the profiles of $J(r)$, in the uppermost layers. First, the internally-generated magnetic field, ${B}_{\rm dyn}$, is usually assumed to be an aligned dipole and orthogonal to ${v}$, but its tilt, and the contributions from higher multipoles, are often non-negligible, as seen both in the Solar system planets (Schubert & Soderlund, 2011) and in dynamo simulations (e.g., Gastine et al. 2014; Wicht et al. 2019b, a; Elias-López et al. 2025b). Secondly, and related to that, an idealized configuration for which ${v}\times\@vec{B}\simeq\@vec{v}\times\@vec{B}_{\rm dyn}$ in eq. (8), implies that either ${B}_{\rm dyn}$ and ${v}$ are perfectly orthogonal, or that the induced field is only a perturbation to the background, ${B}_{\rm ind}\ll\@vec{B}_{\rm dyn}$. The latter approximation breaks down at sufficiently high temperatures ($T\gtrsim 1500$ K for reasonable values of $B_{\rm dyn}\sim 10$ G), due to the increasing magnetic Reynolds number $Rm>$ 1 (Dietrich et al., 2022). Therefore, on one side ${B}$ becomes more complex, and, at the same time, the flow circulation patterns, ${v}$, are distorted by the magnetic field configuration (Batygin et al., 2013; Rogers & Showman, 2014; Rogers & Komacek, 2014). Third, the value of $\sigma$, which is highly sensitive on the local temperature, shows important latitudinal and longitudinal variations (Rogers & McElwaine, 2017) due to the thermal gradients caused by the tidal locking and the impossibility of having a perfectly thermal redistribution of the global circulation. Moreover, the radial profiles of $\sigma$ are greatly affected by the local composition and opacity, as shown for instance by the important differences in the $\sigma(r)$ profiles derived from using different $p(T)$ in modelling a given planet (Kumar et al., 2021). Finally, the 1D evolutionary calculations typically don’t include the uppermost layers ($\lesssim$ mbar), where other effects like dayside photoionization (Koskinen et al., 2014), ambipolar diffusion (Savel et al., 2024; Soriano-Guerrero et al., 2025), current leakage to the magnetosphere can take place and affect the atmospheric electrodynamics, and thermo-resistive instability, which potentially arises from magnetic drag in the dayside and can lead to a thermal runaway as the conductivity increases (Menou, 2012b; Hardy et al., 2022, 2023, 2025).
@@
Also for these reasons, and in contrast with previous pioneering efforts (Batygin & Stevenson, 2010; Batygin et al., 2011; Wu & Lithwick, 2013), here we do not aim at including a mathematically detailed solution for the atmospheric region in the 1D evolutionary code. Instead, below a certain pressure $p_{\rm atm}$ where the term ${v}\times\@vec{B}$ is non-negligible, we estimate the order of magnitude of the sphere-averaged current density profile by the following effective parametrization:
@@ 4.2.2 Induced currents in deeper regions
In deeper regions, $p\geq p_{\rm atm}$, wind intensity decays quickly, so that we neglect the ${v}\times\@vec{B}$ term in eq. (9), and find a solution for $\Phi$. Moreover, no strong temperature differences are expected, since the $p(T)$ profile converges to the internal adiabat: 1D, spherical symmetry calculations are then appropriate and the conductivity gradient is purely radial. Therefore, eq. (9), for each multipole $\Phi_{lm}$, reduces to:
@@
where the primes denote radial derivatives, and we made use of the spherical harmonic identities for the angular part of the Laplacian operator, to get the $l(l+1)/r^{2}$ term. In general, eq. (12) is solved numerically for a given profile of $\sigma$, see e.g. Batygin & Stevenson (2010); Batygin et al. (2011); Wu & Lithwick (2013) for specific modes.
Analytical solutions exist if we parametrize the conductivity with a power law, $\sigma\sim r^{-\beta}$, so that:
@@
We will stick only to the positive branch, since we perform our calculation down to the envelope-core interface, and the negative branch would imply diverging solutions toward the centre of the planet, as $r\rightarrow 0$.
Therefore, we have the sphere-average radial profile |${\nabla}\Phi|_{s}(r)\sim\sum_{l}j_{l}r^{\alpha(\beta,l)-1}$, where $j_{l}$ represent the relative weights of the contributions from each multipole degree $l$ (summing all the orders $m\in[-l,l]$ of a given $l$, and averaging over the sphere), and are taken as a free set of effective parameters of the model, rather than computed for a precise configuration of ${v}(r,\theta,\phi)$ and ${B}(r,\theta,\phi)$ (like in e.g. Batygin & Stevenson 2010; Batygin et al. 2011; Knierim et al. 2022). The current density profile is then given by
@@
$Q_{j}\sim j_{l}^{2}r^{2\alpha(\beta,l)-\beta-2}\sim j_{l}^{2}r^{\sqrt{(\beta-1)^{2}+4l(l+1)}-3}\,.$ (16)
For the simplest configuration employed by Batygin & Stevenson (2010), $l=2$ is the only mode, $j_{2}\neq 0$. In the simplified case $\sigma^{\prime}=\beta=0$, eq. (12) reduces to the Laplace equation, with solutions $\Phi_{lm}\sim r^{l}$. Moreover, in the $\beta=0$, $l=2$ case, one recovers the simple case $\alpha=2$, $J\sim r$, and $Q_{j}\sim r^{2}$, assumed by Ginzburg & Sari (2016). Any other value in the range $\beta>0$ (which is expected everywhere except close to the irradiation region, see Fig. 2), $j_{l>2}\neq 0$ (complexity in the wind and magnetic field geometries) will give higher values of $\alpha$ in the positive branch, since it is a monotonic function of both $l$ and $\beta$. This means that the profile of $J$ will increase more steeply with $r$ for more complex configurations of the wind and/or the dynamo magnetic field (high $l$), and for steeper conductivity values (high $\beta$). In the limit of large values of $\beta\gg l$ (steep radial gradient of $\sigma$), $\alpha\rightarrow\left[\beta-1+\frac{l(l+1)}{\beta-1}\right]$, so that $J\sim r^{-2+l(l+1)/(\beta-1)}$ and $Q_{j}\sim r^{\beta-4+2l(l+1)/(\beta-1)}$, i.e., a steeply decreasing Ohmic dissipation profile with decreasing radial distance.
@@
This semi-analytical approach allows us to easily take into account all $\Phi_{lm}$, and to test any combination of $j_{l}$, i.e., of the relative contribution of each multipole degree $l$ to |${\nabla}\Phi|_{s}$, given by wind and magnetic field configurations in the atmospheric layers. In order to consider a realistic set of relative weights $j_{l}$, we need to consider the geometries of the deep-seated magnetic field and the velocity field.
@@
Secondly, the wind geometry arguably shows stronger jets in the equatorial regions, with possible counter-jets at higher latitudes. Usually, an approximate equatorial symmetry is expected. If one takes the simplest possible case like in Batygin & Stevenson (2010), $v_{\phi}\sim\sin\theta$ and an aligned, dipolar field, one has $j_{l}\neq 0$ for $l=2$, which can be simplistically seen as a combination of the $l_{w}=1$ and $l_{b}=1$ contributions from the wind and magnetic fields, respectively. Generalizing it to a given wind contribution $l_{w}$, and a set of magnetic contributions $b_{l}(R_{\rm dyn})$, the weights of the currents of eq. (15) can be parametrized by:
@@
This expression is an effective, practical proxy to the more rigorous and case-to-case variable decomposition of the atmospheric $\sigma(\@vec{v}\times\@vec{B})$ term, which depends on the specific, detailed profiles of $\sigma$, velocity and magnetic fields. Below, we will consider cases ranging from a pure dipole (only $b_{1}\neq 0$), to a flat, continuum distribution of magnetic weights ($b_{l}(R_{\rm dyn})$ constant) up to a maximum multipole $l_{max}$.
Figure 4: Current reconstruction for the two models with heating considered in Fig. 2. Top: evolution of the best-fitting power law indexes $\beta$ (black, indicating the steepness of the conductivity profile $\sigma(r)\sim r^{-\beta}$) at shallow ($p\sim 10^{2}$ bar, solid), intermediate ($p\sim 10^{4}$ bar, dashed) and deep ($p\sim 10^{6}-10^{7}$ bar, dots) layers, and the corresponding local indexes $\alpha$ of the induced current profile, eq. (14), for $l=2$ (cyan) and $l=10$ (green). Middle: radial profiles of the induced currents for the same model, at ages of $t\simeq 0.4$ (dots), 1.1 (dashes), and 5 Gyr (solid lines), for a purely dipolar case, (only $j_{2}\neq 0$, blue and red) or multipolar, with a flat magnetic spectrum at dynamo surface from $l=2$ to $l=10$ (cyan and golden). Bottom: corresponding Ohmic heating rate profiles.
In summary, at each time step, by performing a series of piecewise power-law fits of the conductivity profiles $\sigma(r)$, we can reconstruct the current density profile, testing a broad range of geometries of the magnetic and wind fields. In the top panel of Fig. 3 we show the reconstructed induced currents for the conductivity profiles of Fig. 2, discussed above, for the purely aligned dipolar case (only $j_{2}\neq 0$). In most of the domain, the conductivity increases inward with $\beta\gg l$, so that the average induced current density has a very slight increase inward, following the $J\sim r^{-2}$ limit derived above. As we go inward into the degeneracy-dominated conductivity, $\beta$ is only slightly positive, so that the current density gradually reaches a maximum and then approaches the limit $J\sim r$. Overall, notice that $J$ varies only by a factor of a few between the exterior and the beginning of the dynamo region.
@@
In Fig. 4 we consider the same two models with heating considered in Fig. 2. In the top panel, we show the evolution of the radial indexes $\beta$ at three different layers for the $T_{\rm eq}=1500$ K model, and the corresponding values of $\alpha$ (eq. 14) for the cases $l=2$ (the simplest, usually considered), and $l=10$ (representative of contributions from high multipoles). In the middle and bottom panels, we show the corresponding solutions for $J$ and $\epsilon_{j}$ for different sets of $j_{l}$ at three ages, for the $T_{\rm eq}=1500$ and 2000 K heated models, and the simplest topology $j_{2}\neq 0$, or a flat distribution at the dynamo, with non-zero, equal contributions $b_{l}$ for $j\in[1,9]$, and $l_{w}=1$ (hence, $j_{l}\neq 0$ for $j\in[2,10]$, eq. 17). The only differences are seen at very deep layers, where, in any case, the dynamo is operating and, therefore, the interaction between the induced and dynamo currents should be formally considered (see discussion below).
The reason for the lack of relevant differences is that, in most of the domain, the solution is determined by the steepness of the conductivity profile, $\beta\gg l$, and only when $\beta\sim l$ the magnetic (and wind) geometry matters. This reinforces and complements what Batygin et al. (2011) found for exact solutions to different specific wind and magnetic geometries: they found only a minor variability in the induced current profile, mostly driven by the wind region precise solution for a given geometry (which here we instead parametrize with $v_{\rm avg}$). Keeping the same normalization, the differences in the integrated Ohmic heat (hence, in the Ohmic efficiency and the overall evolution) are negligible. We will not explore this further and, hereafter, all models include only contribution from $j_{2}$.

### v2

@@ 4.2 Atmospherically induced current
Our aim is to estimate the spherically averaged radial profile of the current density, $J(r,t):=<|\@vec{J}|>_{\theta,\phi}(r,t)$, where ${J}=(\@vec{\nabla}\times\@vec{B})/\mu_{0}$ and $\mu_{0}$ is the magnetic permeability, in order to evaluate the radial profile of the local Joule dissipation rate, $Q_{\rm j}(r,t)=J(r,t)^{2}/\sigma(r,t)$, in the 1D code. As in the first works on Ohmic dissipation in HJs (Batygin & Stevenson, 2010; Batygin et al., 2011), we quantified the electrical currents by imposing a stationary solution to the induction equation given by the balance between the advection (itself given in terms of the thermal wind ${v}$) and the resistive term (neglecting the Hall and ambipolar non-ideal terms, see Perna et al. 2010a; Koskinen et al. 2014; Soriano-Guerrero et al. 2025):
$\frac{d\@vec{B}_{\rm ind}}{dt}=\@vec{\nabla}\times\left(\@vec{v}\times\@vec{B}-\frac{\@vec{J}}{\sigma}\right)=0\leavevmode\nobreak\ .$ (7)
Here, the magnetic field, ${B}$, includes the dynamo-generated background field, ${B}_{\rm dyn}$, and the atmospherically induced field itself, ${B}_{\rm ind}$. Following e.g. Batygin & Stevenson (2010), at equilibrium, the solution of the induced current density could then be defined by considering the electrical potential, $\Phi$,
@@
and imposing the continuity equation, ${J}=0$, so that
@@
We then expanded $\Phi$ in spherical harmonics, $\Phi=\Sigma_{lm}\Phi_{lm}(r)Y_{l}^{m}(\theta,\phi)$, and solved for the radial dependence of each multipole, $\Phi_{lm}(r)$, for a given prescription of ${v}$, ${B}$, and $\sigma$ — which in turn depends on $T$ and density, i.e. on $p(T)$. Batygin & Stevenson (2010), Batygin et al. (2011), and related works solve such an elliptic equation, considering first the uppermost regions where ${v}\neq 0$ and assuming axial symmetry (i.e. $\Phi_{lm}=0$ for any $m\neq 0$) and simplified geometries for ${v}(r,\theta)$, ${B}(r,\theta)$ for a given profile $\sigma(r)$. The simplest configuration is a purely zonal wind $v_{\phi}\sim\sin\theta$ ($\theta\in[0,\pi]$ being the co-latitude) and a dipolar, aligned field, which means that the only possible induced current multipole is the quadrupolar one, $\Phi_{20}\neq 0$. Here we propose new families of semi-analytical solutions for the deep layers where ${v}\sim 0$, treating the complex atmospheric layers in an effective, parametrised way and allowing for any generic profile $\sigma(r)$.
Figure 2: Electrical conductivity, $\sigma$, as a function of pressure for different planetary models. We show ages of $t\simeq 0.4$ (dots), 1.1 (dashes), and 5 Gyr (solid lines) for a planet with $M=1\leavevmode\nobreak\ M_{\rm j}$, showing two cases, $T_{\rm eq}=1500,2000$ K, without (black, grey) and with Ohmic heating, parameter $v_{\rm avg}=3000$ and 356 m/s, respectively (blue, red). We also show (brown and gold lines), for comparison, the simplified $K^{+}$-only thermal ionisation formula, which neglects pressure ionisation, for the cases without Ohmic heating.
Figure 3: Profiles as a function of pressure of the induced currents, $J$ (top), and Ohmic specific (middle) and cumulative (bottom) heat rate. We show the same heated cases and ages as in Fig. 2. The thick lines in the bottom panel identify the convective regions at each age of each model.
@@ 4.2.1 Treatment of the induction in the atmospheric region
There are several factors that introduce uncertainties and, likely, planet-to-planet variations in the profiles of $J(r)$, in the uppermost layers. First, the internally generated magnetic field, ${B}_{\rm dyn}$, is usually assumed to be an aligned dipole and orthogonal to ${v}$, but its tilt, and the contributions from higher multipoles, are often non-negligible, as seen both in the Solar system planets (Schubert & Soderlund, 2011) and in dynamo simulations (e.g. Gastine et al. 2014; Wicht et al. 2019b, a; Elias-López et al. 2025b). Secondly, and related to that, an idealised configuration for which ${v}\times\@vec{B}\simeq\@vec{v}\times\@vec{B}_{\rm dyn}$ in eq. (8), implies that either ${B}_{\rm dyn}$ and ${v}$ are perfectly orthogonal, or that the induced field is only a perturbation to the background, ${B}_{\rm ind}\ll\@vec{B}_{\rm dyn}$. The latter approximation breaks down at sufficiently high temperatures ($T\gtrsim 1500$ K for reasonable values of $B_{\rm dyn}\sim 10$ G), due to the increasing magnetic Reynolds number $Rm>$ 1 (Dietrich et al., 2022). Therefore, on one side ${B}$ becomes more complex, and, at the same time, the flow circulation patterns, ${v}$, are distorted by the magnetic field configuration (Batygin et al., 2013; Rogers & Showman, 2014; Rogers & Komacek, 2014). Third, the value of $\sigma$, which is highly sensitive on the local temperature, shows important latitudinal and longitudinal variations (Rogers & McElwaine, 2017) due to the thermal gradients caused by the tidal locking and the impossibility of having a perfectly thermal redistribution of the global circulation. Moreover, the radial profiles of $\sigma$ are greatly affected by the local composition and opacity, as shown for instance by the important differences in the $\sigma(r)$ profiles derived from using different $p(T)$ in modelling a given planet (Kumar et al., 2021). Finally, the 1D evolutionary calculations typically do not include the uppermost layers ($\lesssim$ mbar), where other effects like dayside photoionisation (Koskinen et al., 2014), ambipolar diffusion (Savel et al., 2024; Soriano-Guerrero et al., 2025), current leakage to the magnetosphere can take place and affect the atmospheric electrodynamics, and thermo-resistive instability, which potentially arises from magnetic drag in the dayside and can lead to a thermal runaway as the conductivity increases (Menou, 2012b; Hardy et al., 2022, 2023, 2025).
@@
Also for these reasons, and in contrast with previous pioneering efforts (Batygin & Stevenson, 2010; Batygin et al., 2011; Wu & Lithwick, 2013), here we do not aim at including a mathematically detailed solution for the atmospheric region in the 1D evolutionary code. Instead, below a certain pressure $p_{\rm atm}$ where the term ${v}\times\@vec{B}$ is non-negligible, we estimate the order of magnitude of the sphere-averaged current density profile by the following effective parametrisation:
@@ 4.2.2 Induced currents in deeper regions
In deeper regions, $p\geq p_{\rm atm}$, wind intensity decays quickly, so that we neglect the ${v}\times\@vec{B}$ term in eq. (9), and find a solution for $\Phi$. Moreover, no strong temperature differences are expected, since the $p(T)$ profile converges to the internal adiabat: 1D, spherical symmetry calculations are then appropriate and the conductivity gradient is purely radial. Therefore, eq. (9) for each multipole, $\Phi_{lm}$, reduces to
@@
where the primes denote radial derivatives, and we made use of the spherical harmonic identities for the angular part of the Laplacian operator to get the $l(l+1)/r^{2}$ term. In general, eq. (12) is solved numerically for a given profile of $\sigma$ (see e.g. Batygin & Stevenson (2010); Batygin et al. (2011); Wu & Lithwick (2013) for specific modes).
Analytical solutions exist if we parametrise the conductivity with a power law, $\sigma\sim r^{-\beta}$, so that
@@
We stick only to the positive branch, since we performed our calculation down to the envelope-core interface, and the negative branch would imply diverging solutions towards the centre of the planet as $r\rightarrow 0$.
Therefore, we obtained the sphere-average radial profile |${\nabla}\Phi|_{s}(r)\sim\sum_{l}j_{l}r^{\alpha(\beta,l)-1}$, where $j_{l}$ represent the relative weights of the contributions from each multipole degree, $l$ (summing all the orders $m\in[-l,l]$ of a given $l$ and averaging over the sphere), and are taken as a free set of effective parameters of the model rather than computed for a precise configuration of ${v}(r,\theta,\phi)$ and ${B}(r,\theta,\phi)$, as in Batygin & Stevenson 2010; Batygin et al. 2011; Knierim et al. 2022, for example. The current density profile is then given by
@@
$Q_{\rm j}\sim j_{l}^{2}r^{2\alpha(\beta,l)-\beta-2}\sim j_{l}^{2}r^{\sqrt{(\beta-1)^{2}+4l(l+1)}-3}\,.$ (16)
For the simplest configuration employed by Batygin & Stevenson (2010), $l=2$ is the only mode, $j_{2}\neq 0$. In the simplified case $\sigma^{\prime}=\beta=0$, eq. (12) reduces to the Laplace equation, with solutions $\Phi_{lm}\sim r^{l}$. Moreover, in the $\beta=0$, $l=2$ case, one recovers the simple case $\alpha=2$, $J\sim r$, and $Q_{\rm j}\sim r^{2}$, assumed by Ginzburg & Sari (2016). Any other value in the range $\beta>0$ (which is expected everywhere except close to the irradiation region, see Fig. 2), $j_{l>2}\neq 0$ (complexity in the wind and magnetic field geometries) will give higher values of $\alpha$ in the positive branch, since it is a monotonic function of both $l$ and $\beta$. This means that the profile of $J$ will increase more steeply with $r$ for more complex configurations of the wind and/or the dynamo magnetic field (high $l$), and for steeper conductivity values (high $\beta$). In the limit of large values of $\beta\gg l$ (steep radial gradient of $\sigma$), $\alpha\rightarrow\left[\beta-1+\frac{l(l+1)}{\beta-1}\right]$, so that $J\sim r^{-2+l(l+1)/(\beta-1)}$ and $Q_{\rm j}\sim r^{\beta-4+2l(l+1)/(\beta-1)}$, i.e. a steeply decreasing Ohmic dissipation profile with decreasing radial distance.
@@
This semi-analytical approach allows us to easily take into account all $\Phi_{lm}$, and to test any combination of $j_{l}$, i.e. of the relative contribution of each multipole degree $l$ to |${\nabla}\Phi|_{s}$, given by wind and magnetic field configurations in the atmospheric layers. In order to consider a realistic set of relative weights $j_{l}$, we need to consider the geometries of the deep-seated magnetic field and the velocity field.
@@
Secondly, the wind geometry arguably shows stronger jets in the equatorial regions, with possible counter-jets at higher latitudes. Usually, an approximate equatorial symmetry is expected. If one takes the simplest possible case like in Batygin & Stevenson (2010), $v_{\phi}\sim\sin\theta$ and an aligned, dipolar field, one has $j_{l}\neq 0$ for $l=2$, which can be simplistically seen as a combination of the $l_{w}=1$ and $l_{b}=1$ contributions from the wind and magnetic fields, respectively. Generalizing it to a given wind contribution, $l_{w}$, and a set of magnetic contributions, $b_{l}(R_{\rm dyn})$, the weights of the currents of eq. (15) can be parametrised by
@@
This expression is an effective, practical proxy to the more rigorous and case-to-case variable decomposition of the atmospheric $\sigma(\@vec{v}\times\@vec{B})$ term, which depends on the specific, detailed profiles of $\sigma$, velocity and magnetic fields. Below, we consider cases ranging from a pure dipole (only $b_{1}\neq 0$), to a flat, continuum distribution of magnetic weights ($b_{l}(R_{\rm dyn})$ constant) up to a maximum multipole $l_{max}$.
Figure 4: Current reconstruction for the two models with heating considered in Fig. 2. Top: Evolution of the best-fitting power-law indexes, $\beta$ (black, indicating the steepness of the conductivity profile $\sigma(r)\sim r^{-\beta}$), at shallow ($p\sim 10^{2}$ bar, solid), intermediate ($p\sim 10^{4}$ bar, dashed), and deep ($p\sim 10^{6}-10^{7}$ bar, dots) layers and the corresponding local indexes, $\alpha$, of the induced current profile, eq. (14), for $l=2$ (cyan) and $l=10$ (green). Middle: Radial profiles of the induced currents for the same model at ages of $t\simeq 0.4$ (dots), 1.1 (dashes), and 5 Gyr (solid lines) for a purely dipolar case, (only $j_{2}\neq 0$, blue and red) or multipolar, with a flat magnetic spectrum at dynamo surface from $l=2$ to $l=10$ (cyan and golden). Bottom: Corresponding Ohmic heating rate profiles.
In summary, at each time step, by performing a series of piecewise power-law fits of the conductivity profiles $\sigma(r)$, we can reconstruct the current density profile, testing a broad range of geometries of the magnetic and wind fields. In the top panel of Fig. 3 we show the reconstructed induced currents for the conductivity profiles of Fig. 2, discussed above, for the purely aligned dipolar case (only $j_{2}\neq 0$). In most of the domain, the conductivity increases inward with $\beta\gg l$, so that the average induced current density has a very slight increase inward, following the $J\sim r^{-2}$ limit derived above. As we go inward into the degeneracy-dominated conductivity, $\beta$ is only slightly positive, so that the current density gradually reaches a maximum and then approaches the limit $J\sim r$. Overall, we noticed that $J$ varies only by a factor of a few between the exterior and the beginning of the dynamo region.
@@
In Fig. 4 we consider the same two models with heating considered in Fig. 2. In the top panel, we show the evolution of the radial indexes $\beta$ at three different layers for the $T_{\rm eq}=1500$ K model, and the corresponding values of $\alpha$ (eq. 14) for the cases $l=2$ (the simplest, usually considered), and $l=10$ (representative of contributions from high multipoles). In the middle and bottom panels, we show the corresponding solutions for $J$ and $\epsilon_{\rm j}$ for different sets of $j_{l}$ at three ages, for the $T_{\rm eq}=1500$ and 2000 K heated models, and the simplest topology $j_{2}\neq 0$, or a flat distribution at the dynamo, with non-zero, equal contributions $b_{l}$ for $j\in[1,9]$, and $l_{w}=1$ (hence, $j_{l}\neq 0$ for $j\in[2,10]$, eq. 17). The only differences are seen at very deep layers, where, in any case, the dynamo is operating, and therefore, the interaction between the induced and dynamo currents should be formally considered (see discussion below).
The reason for the lack of relevant differences is that, in most of the domain, the solution is determined by the steepness of the conductivity profile, $\beta\gg l$, and only when $\beta\sim l$ the magnetic (and wind) geometry matters. This reinforces and complements what Batygin et al. (2011) found for exact solutions to different specific wind and magnetic geometries: they found only a minor variability in the induced current profile, mostly driven by the wind region precise solution for a given geometry (which here we instead parametrise with $v_{\rm avg}$). Keeping the same normalisation, the differences in the integrated Ohmic heat (hence, in the Ohmic efficiency and the overall evolution) are negligible. We do not explore this further, and hereafter, all models include only a contribution from $j_{2}$.